# Quokka: Accelerating Program Verification with LLMs via Invariant Synthesis

### v2

@@ 3.2 Verifier-Based Decision Procedure
This theorem establishes that whenever the calculus derives a conclusive outcome, that outcome is correct. The inconclusive case $\mathbf{U}$ is deliberately conservative: it makes no claim about the truth or falsity of the property and safely captures verifier incompleteness or timeouts.
3.3 Implementation
We describe the implementation of our verifier-based evaluation framework. Given a program $P$ and target property $p^{\star}$, the system evaluates the generated invariants according to the decision procedure. Each invariant $q=\langle\psi,\ell\rangle$ consists of a program location $\ell$ and predicate $\psi$.
Syntactic Validation.
Parallel Verifier Queries. For each candidate $q$, the procedure issues two verifier queries, namely $d_{a}=V(P,\varnothing,q)$ to check whether $q$ is an invariant and $d_{b}=V(P,\{q\},p^{\star})$ to check whether the target holds under the assumption $q$. These queries are executed in parallel in our implementation, which reduces latency and enables speedup when the proposed invariant is useful for verification. The final outcome is then derived exactly according to the decision calculus.
3.4 Supervised Fine-Tuning and Best-of-N Sampling
Below, we discuss how we construct our dataset for fine-tuning and the way we perform Best-of-N sampling.
Synthetic Dataset Generation.
To construct the synthetic dataset, we prompt GPT-4o using the template in Appendix A.1. The template takes three seed programs as examples and instructs the model to synthesize a new C program that is compilable and contains both loops and assertions. To obtain a diverse and large pool of candidates, we repeatedly invoke the model with different seed programs. To avoid data leakage, these seed programs are randomly drawn from the SV-COMP pool  (Beyer and Strejček, 2025) that is disjoint from our evaluation set. Although the prompt requests compilable programs with loops and assertions, the LLM-generated programs may fail to compile, include assertions that do not hold, and contain multiple assertions. For programs with multiple assertions, we split them into separate instances, each retaining only a single assertion while preserving all loop structures (ensuring at least one loop per instance). We then run UAutomizer on every program and discard any instance that is non-compilable or whose assertion is invalid. This filtering step ensures the quality of the dataset, resulting in 3589 synthetic programs.
Extract Invariants Generated from UAutomizer.
When running UAutomizer to prove the assertions in the synthetic programs, the tool also emits the invariants it discovers. From its output, we extract the loop invariants. Each extracted invariant includes its program location and its predicate. Although each program contains exactly one assertion, it may include multiple loops, so a single program can yield multiple loop invariants, all associated with the same assertion. We pair each program with each of its corresponding loop invariants to form our training dataset. We show an example invariant generated from UAutomizer in Appendix A.2. With the ground-truth invariants, we perform supervised fine-tuning using LoRA (Hu et al., 2022).
Best-of-N Sampling.
Best-of-N sampling is an inference-time strategy in which multiple candidate programs are generated, and the most effective one is selected, a technique shown to improve performance on code-generation tasks (Ehrlich et al., 2025). In our setting, the best candidate is the invariant that yields the largest speedup, i.e., the one whose decision procedure finishes earliest. As described in Section˜3.2, evaluating a single candidate requires two verifier queries issued in parallel; therefore, Best-of-N sampling evaluates 2N verifier queries concurrently.
4 Experimental Setup
Dataset.
We construct our evaluation benchmark from SV-COMP (Beyer and Strejček, 2025), a standard competition in software verification, focusing on problems that require loop invariant synthesis. We collect a pool of 866 instances and run the state-of-the-art non-LLM verifier UAutomizer (Schüssele et al., 2024) with a 600-second timeout to record per-instance solving times. The process is fully automated, with no manual filtering or cherry-picking. We use this timeout to enable meaningful measurement of verification speedups. In contrast to prior work (Wu et al., 2024b), which evaluates only instances that the baseline cannot solve within this time budget, our approach requires knowing the baseline runtime to quantify relative improvements. Restricting to instances whose baseline runtime is observable allows us to report concrete speedup factors, rather than only counts of additionally solved problems. To the best of our knowledge, this dataset is the largest used in prior evaluations of LLM-based verifiers. Also, the programs in our dataset are significantly longer, with an average of 53 lines of code, compared to only 22 for LaM4Inv (Wu et al., 2024a), 23 for LEMUR (Wu et al., 2024b), and 27 for Loopy (Kamath et al., 2023). Manual inspection shows that our dataset contains features such as multiple loops, functions, arrays, and pointers, which are largely absent from prior datasets. This makes Quokka a more challenging and realistic benchmark that better distinguishes solver performance.
For training data, we construct our training data with synthetic data generation filtered by the solvers, as described in Section˜3.4. We make sure that the seed programs used for synthetic data generation are disjoint from our evaluation dataset, mitigating the leakage issue. The statistics of our dataset are shown in Table˜1.
Dataset Avg. #Lines #Instances
Evaluation 53 866
Training 42 3589
Table 1: Dataset statistics.
Model # Correct Invariants # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
Qwen2.5-7B 419 34 26 23 20 18
o3 550 38 33 30 24 23
Llama-3.1-8B 342 43 35 32 29 27
claude-opus-4.1 487 56 45 38 35 31
Qwen2.5-72B 500 63 49 45 37 33
claude-opus-4.5 689 69 53 46 39 34
gpt-5.1 694 56 46 41 40 37
claude-sonnet-4 620 68 58 51 45 41
gpt-5.2 710 81 65 60 56 51
Table 2: Performance of different LLMs when using Quokka. We report the number of instances with verified-correct invariants and the number of instances achieving speedups of at least $1.2\times$, $1.4\times$, $1.6\times$, $1.8\times$, and $2.0\times$ over UAutomizer.
Metrics.
We evaluate LLMs along two dimensions: (1) the correctness of the generated invariants and (2) the performance improvements they provide. Correctness is determined by the solver using a timeout set to $1.2\times$ the original solving time for each problem. This slack allows the solver additional time to assess the correctness of model-generated invariants, reducing inconclusive outcomes caused by timeouts. All speedup-related measurements are reported relative to the original solver, which serves as the baseline. We report speedup using the metric of $\mathtt{\#Fast}_{p}$, defined as the number of instances for which verification completes at least $p\times$ faster than the baseline, with $p\in\{1.2,1.4,\ldots\}$. Additionally, we also report the number of instances solved under varying timeout budgets.
Choice of Solvers.
We use the state-of-the-art non-LLM-based solver UAutomizer (Schüssele et al., 2024) as our base solver. Our methodology requires that the solver can verify programs without externally-provided invariants, enabling us to measure speedup relative to a functional baseline. While prior work has explored other solvers such as Frama-C (Cuoq et al., 2012) and ESBMC (Gadelha et al., 2018), we select UAutomizer for two key reasons. First, UAutomizer represents the current state-of-the-art in automated verification, outperforming ESBMC on standard benchmarks in SV-COMP. Second, unlike Frama-C, which requires externally-provided invariants and cannot proceed without them, UAutomizer has built-in invariant synthesis capabilities that allow it to verify programs autonomously. To demonstrate the generality of Quokka’s methodology across different solvers, we also present experiments with ESBMC in Appendix A.5, where we show that Quokka successfully accelerates ESBMC as well.
@@ 4 Experimental Setup
5 Results
5.1 Results of LLMs
Table˜2 presents the performance of different LLMs when using Quokka. The results are sorted by the number of instances achieving at least $2.0\times$ speedup (# Fast2.0), which ranges from 18 instances for Qwen2.5-7B to 51 instances for gpt-5.2. Overall, the table demonstrates substantial variation in both correctness and speedup performance across different models, with larger and more recent models generally achieving better results.
Effectiveness of Quokka.
Our approach demonstrates the speedup capabilities across all evaluated models. Even smaller models like Qwen2.5-7B achieve speedups on 34 instances at the $1.2\times$ threshold, while the best-performing model gpt-5.2 accelerates 81 instances at this level and achieves $2.0\times$ or greater speedup on 51 instances. These results demonstrate that Quokka can effectively leverage LLMs to generate invariants that meaningfully accelerate verification. The consistent speedup gains across different model families, from 7B parameter models to frontier models, and the progressive improvement with model scale underscore the effectiveness of our approach.
Correctness vs. Performance Gap.
A striking observation is the substantial gap between generating correct invariants and achieving meaningful speedups. For example, gpt-5.2 produces 710 correct invariants but achieves $2.0\times$ speedup on only 51 instances, while claude-opus-4.5 generates 689 correct invariants yet delivers $2.0\times$ speedup on merely 34 instances. This gap highlights that correctness alone is insufficient: invariants must also be strong enough to meaningfully accelerate verification, requiring not only logical correctness but also strategic insight into which invariants will most effectively guide the solver.
Model Inference Time.
We include model inference time in our evaluation to provide a fair and realistic assessment of end-to-end performance. As documented in Appendix A.4, smaller models such as Llama-3.1-8B and Qwen2.5-7B exhibit negligible generation times (0.3–0.4 seconds), while most models complete inference within 3 seconds. The o3 model has the highest overhead at 8.1 seconds on average. Importantly, unlike LOOPY (Kamath et al., 2023), which invokes the LLM up to 15 times per program and thus spends a substantial portion of its time budget on API calls, our approach requires only a single LLM invocation per program. Consequently, LLM inference accounts for only a small fraction of total runtime. Furthermore, our evaluation dataset contains sufficiently challenging programs (with solving times up to 600 seconds) such that solver time naturally dominates inference overhead, making the comparison meaningful even when inference latency is included.
Prompting-Based Methods.
We investigate the impact of different prompting strategies on invariant generation quality using gpt-5.2. As detailed in Appendix A.6, we evaluate 0-shot, 2-shot, 4-shot, and chain-of-thought (CoT) prompting. Few-shot prompting does not consistently improve performance, while CoT prompting yields the weakest results, likely due to increased generation time without proportional quality gains. These findings indicate that simple zero-shot prompting is most effective for invariant synthesis.
5.2 Comparison with Prior LLM-based Verifiers
Method # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
LaM4Inv 1 1 1 1 1
LOOPY 8 7 5 3 1
LEMUR-original 19 16 13 12 10
LEMUR-gpt-5.2 39 30 28 25 22
Quokka-gpt-5.2 81 65 60 56 51
Table 3: Comparison of speedup performance across different methods. We report the number of instances achieving speedups of at least $1.2\times$, $1.4\times$, $1.6\times$, $1.8\times$, and $2.0\times$ over UAutomizer.
Table˜3 compares the speedup performance of Quokka against prior LLM-based verifiers across different speedup thresholds, while Figure˜2 shows the number of solved instances over time for each method.
Quokka Outperforms Prior LLM-based Verifiers.
Quokka with gpt-5.2 achieves state-of-the-art results across all speedup metrics. Among prior work, LEMUR (Wu et al., 2024b) represents the strongest baseline. To ensure a fair comparison, we enhance LEMUR by replacing its original model (gpt-4) with gpt-5.2, creating LEMUR-gpt-5.2. Even with this enhancement, Quokka-gpt-5.2 significantly outperforms LEMUR-gpt-5.2: at the $1.2\times$ speedup threshold, Quokka-gpt-5.2 accelerates 81 instances versus 39 for LEMUR-gpt-5.2 and 19 for LEMUR-original. At $2.0\times$ speedup, Quokka-gpt-5.2 achieves 51 instances compared to 22 for LEMUR-gpt-5.2 and 10 for LEMUR-original. LaM4Inv (Wu et al., 2024a) and LOOPY (Kamath et al., 2023) demonstrate minimal speedup capabilities, achieving at most 1 and 8 instances respectively at the $1.2\times$ level. These results demonstrate that Quokka establishes a new state-of-the-art for LLM-based verification acceleration.
Differences in Base Solvers.
Different LLM-based verification frameworks are built on different base solvers: Quokka and LEMUR use UAutomizer, LOOPY uses Frama-C (Cuoq et al., 2012), and LaM4Inv uses ESBMC (Gadelha et al., 2018). The choice of base solver significantly impacts the performance, as stronger solvers provide a better foundation for speedup. Since UAutomizer represents the current state-of-the-art solver, we believe future work should prioritize developing LLM-based verifiers atop state-of-the-art solvers such as UAutomizer. For a direct comparison between UAutomizer and prior LLM-based solvers, we also refer readers to Appendix A.3, where we evaluate UAutomizer on the dataset released by each of the prior works.
Algorithmic Simplicity and Parallelism.
Prior work predominantly adopts complex, sequential, and iterative processes with highly customized algorithms to handle LLM-generated invariants. In contrast, Quokka employs a simple, verifier-based, and parallel approach that directly leverages the base solver’s capabilities.
LaM4Inv uses a “query-filter-reassemble” strategy that filters and combines model-generated predicates through a complex algorithm, rather than using them directly. LOOPY employs a highly-customized Houdini algorithm with iterative fixing to filter candidates. LEMUR is closest to our approach, also directly using the base solver to verify the correctness of invariants, but follows a purely sequential decision process with backtracking. In contrast, Quokka directly uses LLM-generated invariants without decomposition or reassembly, and exploits parallelism by issuing the verification queries in parallel. This design reduces implementation complexity while enabling more efficient hardware utilization, contributing to the superior speedup performance demonstrated in our experiments.
Model # Incorrect Invariants # Goal Timeout # Invariant Timeout # Both Timeout
gpt-5.1 78 96 245 110
claude-sonnet-4 110 160 209 145
gpt-5.2 56 139 230 121
Table 4: Failure mode breakdown for the top three models on Quokka. Incorrect Invariants: the invariant is refuted by the solver. Goal Timeout: the invariant is verified, but proving the assertion under it times out. Invariant Timeout: verifying the invariant itself times out. Both Timeout: both verification queries time out.
Model # Correct Invariants # Fast1.0 # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
Qwen2.5-7B (base) 419 116 34 26 23 20 18
Qwen2.5-7B (finetuned) 431 141 39 27 26 24 20
5.3 Failure Mode Analysis
To understand why many invariants fail to accelerate verification in Quokka, we analyze failure modes for the top three models. We categorize failures into four types: 1) Incorrect Invariants: the invariant is refuted by the solver; 2) Goal Timeout: the invariant is verified, but proving the assertion under it times out; 3) Invariant Timeout: verifying the invariant itself times out; 4) Both Timeout: both verification queries time out. The results in Table˜4 reveal that while some failures stem from incorrect invariants, the dominant issue is timeout-related. This suggests that current models lack an understanding of what makes verification queries tractable for symbolic solvers. Future work should explore training reward models or other mechanisms to help models internalize solver difficulty, enabling them to propose invariants that genuinely simplify verification by decomposing hard problems into easier subproblems rather than inadvertently generating verification conditions that are as difficult or harder to solve than the original proof goal.
Figure 3: Effect of Best-of-N sampling on the number of instances under different Fastp settings.
5.4 Results of Fine-Tuning and Best-of-N Sampling
Modest Improvement from Supervised Fine-Tuning.
We perform supervised fine-tuning using LoRA (Hu et al., 2022) on Qwen2.5-7B for 3 epochs. As shown in Table˜5, fine-tuning yields modest improvements: correct invariants increase from 419 to 431, and instances achieving meaningful speedups improve across all thresholds. For example, Fast1.0 improves from 116 to 141, and Fast1.2 improves from 34 to 39. These results suggest that domain-specific fine-tuning can improve both correctness and speedup, though gains remain moderate.
Benfits of Best-of-N Sampling.
We evaluate Best-of-N sampling by generating multiple invariant candidates from gpt-5.2 at temperature 0.7 and selecting the one yielding the fastest verification time. As shown in Figure˜3, increasing N from 1 to 4 consistently improves speedup across all Fastp thresholds. However, performance plateaus beyond N=8 due to hardware constraints: our 20-core CPU must verify 2N queries in parallel, and when N exceeds 10, resource contention may negate sampling benefits. Memory limits can further constrain scalability. While Best-of-N sampling is effective, its benefits are hardware-bounded, though modest values of N yield meaningful improvements.
6 Conclusion
This work introduced Quokka, a simple yet effective framework for LLM-based invariant synthesis that accelerates program verification. Our approach employs a principled verification procedure that decomposes the original proof goal into two independent subproblems and launches verifier queries in parallel, providing sound evaluation while achieving state-of-the-art speedup results. Using a benchmark of 866 instances, we evaluated 9 state-of-the-art LLMs across multiple model families and compared against prior LLM-based verifiers. The results demonstrate that Quokka consistently outperforms all existing LLM-based verifiers, achieving speedups of at least 1.2× on 81 instances compared to 39 instances for the previous best approach. We further demonstrated that supervised fine-tuning and Best-of-N sampling can yield measurable improvements in accelerating verification. Our work establishes a new state-of-the-art for LLM-based program verification and provides a principled foundation for future research in this direction.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Appendix
A.1 Prompt Template Used for Training Dataset Generation
@@
Table˜A1 shows the prompt template used for synthesizing training programs from seed programs.
@@
A.2 An Example from the Fine-Tuning Dataset
@@
Figure˜A1 shows an example from the fine-tuning dataset with the program to the UAutomizer and the generated loop invariant.
@@
A.3 Results of UAutomizer on Prior Datasets
@@
A.4 Inference Time of Different Models
Table˜A3 presents the average inference time per instance for different LLMs when generating invariants. As expected, smaller open-source models such as Llama-3.1-8B and Qwen2.5-7B exhibit the fastest generation times, averaging around 0.3–0.4 seconds per instance. Mid-sized models like Qwen2.5-72B and the GPT-5 series require approximately 0.6–1.4 seconds. The Claude models take longer at around 1.6–2.7 seconds per instance. The o3 model has the highest inference overhead, averaging over 8 seconds per instance. These results highlight the trade-off between model capability and inference cost, which must be considered when evaluating end-to-end verification performance.
@@ A.4 Inference Time of Different Models
Model # Correct Invariants # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
gpt-5.1 399 10 9 9 9 9
gpt-5.2 455 17 11 10 9 9
Table A4: Performance of different LLMs when using ESBMC as the base verifier. We report the number of instances with verified-correct invariants and the number of instances achieving speedups of at least $1.2\times$, $1.4\times$, $1.6\times$, $1.8\times$, and $2.0\times$ over ESBMC.
A.5 Results for ESBMC as the Base Solver
To demonstrate the generality of Quokka’s methodology, we conduct additional experiments using ESBMC (Gadelha et al., 2018) as the base solver instead of UAutomizer. As shown in Table˜A4, Quokka successfully accelerates ESBMC as well, with both gpt-5.1 and gpt-5.2 generating hundreds of correct invariants and achieving speedups on multiple instances. For example, gpt-5.2 produces 455 correct invariants and achieves speedups of at least $1.2\times$ on 17 instances. These results confirm that Quokka is not limited to a specific solver and can be applied to accelerate different verification tools.
A.6 Results for Prompting-Based Methods
Model # Correct Invariants # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
gpt-5.2 (CoT) 639 41 32 26 22 20
gpt-5.2 (4-shot) 756 72 62 53 47 43
gpt-5.2 (2-shot) 764 83 67 58 53 49
gpt-5.2 710 81 65 60 56 51
@@ A.6 Results for Prompting-Based Methods
We investigate the impact of different prompting strategies on invariant generation quality using gpt-5.2. As shown in Table˜A5, we compare chain-of-thought (CoT) prompting, few-shot prompting with 2 and 4 examples, and our default zero-shot approach. The results show that few-shot prompting does not necessarily improve performance and may introduce noise or distract the model from the task at hand. Chain-of-thought prompting yields the weakest results across all metrics, with only 639 correct invariants and 41 instances achieving $1.2\times$ speedup. This suggests that explicit reasoning steps may increase generation time without proportional benefits to invariant quality, thereby reducing overall speedup. These findings indicate that prompting-based strategy can hardly improve the task of invariant synthesis.

### v3

@@ A.1 Soundness Proof
A.2 Additional Details on Experiment Setup
@@
A.3 Prompt Template Used for Training Dataset Generation
@@
Table A1 shows the prompt template used for synthesizing training programs from seed programs.
@@
A.4 Prompt Template Used for Invariant Generation
Table A2 shows the prompt template used to prompt the LLM to generate a loop invariant for a given C program.
System prompt:
Given a C program with line numbers, you need to identify the loop invariant to be added at one of the provided points in the program. A loop invariant is an expression that is true before, during, and after the execution of a loop.
You will be provided the set of points in the program where loop invariants can be added. Each point will be specified as after line X.
Choose one of the points to insert a loop invariant. Respond with exactly this format:
"After line X, insert assume(condition);"
Where:
- X refers to the provided line number where the loop invariant needs to be added
- condition is the loop invariant condition you want to insert. The loop invariant should not modify any variable (no ++, --, +=, -=, =).
- only output one loop invariant
User prompt:
Program:
{PROGRAM}
Choose one of the points to insert a correct loop invariant at in the program:
{POINTS}
Please analyze the provided C program and output only where to insert a loop invariant in the specified format.
Table A2: Prompt template used to generate a loop invariant for a given program. The placeholder {PROGRAM} denotes the input C program with line numbers, and {POINTS} denotes the candidate locations where the invariant may be inserted.
A.5 An Example from the Fine-Tuning Dataset
@@
Figure A1 shows an example from the fine-tuning dataset with the program to the UAutomizer and the generated loop invariant.
@@
A.6 Results of UAutomizer on Prior Datasets
@@
A.7 Inference Time of Different Models
Table A4 presents the average inference time per instance for different LLMs when generating invariants. As expected, smaller open-source models such as Llama-3.1-8B and Qwen2.5-7B exhibit the fastest generation times, averaging around 0.3–0.4 seconds per instance. Mid-sized models like Qwen2.5-72B and the GPT-5 series require approximately 0.6–1.4 seconds. The Claude models take longer at around 1.6–2.7 seconds per instance. The o3 model has the highest inference overhead, averaging over 8 seconds per instance. These results highlight the trade-off between model capability and inference cost, which must be considered when evaluating end-to-end verification performance.
@@ A.7 Inference Time of Different Models
Model # Correct Invariants 60s 500s
$\Delta$ # Solved $\overline{T}$ $\Delta$ # Solved $\overline{T}$
gpt-5.1 399 31 394 34.8 1 399 272.9
gpt-5.2 455 30 450 31.2 1 455 240.9
Table A5: Performance of different LLMs when using ESBMC as the base verifier.
A.8 Results for ESBMC as the Base Solver
To demonstrate the generality of Quokka’s methodology, we conduct additional experiments using ESBMC (Gadelha et al., 2018) as the base solver instead of UAutomizer. As shown in Table A5, Quokka successfully accelerates ESBMC as well, with both gpt-5.1 and gpt-5.2 generating hundreds of correct invariants and solving additional instances. These results confirm that Quokka is not limited to a specific solver and can be applied to accelerate different verification tools.
A.9 Results for Prompting-Based Methods
Model # Correct Invariants 60s 500s
# $\Delta$ # Solved $\overline{T}$ # $\Delta$ # Solved $\overline{T}$
gpt-5.2 (CoT) 639 18 609 33.4 2 638 153.5
gpt-5.2 710 21 681 22.2 1 710 105.1
gpt-5.2 (4-shot) 756 20 721 20.1 0 754 81.1
gpt-5.2 (2-shot) 764 20 730 19.4 2 763 76.0
@@ A.9 Results for Prompting-Based Methods
We investigate the impact of different prompting strategies on invariant generation quality using gpt-5.2. As shown in Table A6, we compare chain-of-thought (CoT) prompting, few-shot prompting with 2 and 4 examples, and our default zero-shot approach. The results show that few-shot prompting can sometimes improve performance (e.g., 2-shot is generally better than 0-shot), but more examples may also introduce noise or distract the model from the task at hand (e.g., 4-shot is worse than 2-shot). Chain-of-thought prompting yields the weakest results across all metrics, with only 639 correct invariants and the fewest solved instances. This suggests that explicit reasoning steps may increase generation time without proportional benefits to invariant quality, thereby reducing overall speedup. These findings indicate that few-shot prompting could improve the task of invariant synthesis.
A.10 Detailed Comparison against Prior Work
Comparison of Base Solver
Different LLM-based verification frameworks are built on different base solvers: Quokka and LEMUR use UAutomizer, LOOPY uses Frama-C (Cuoq et al., 2012), and LaM4Inv uses ESBMC (Gadelha et al., 2018). The choice of base solver significantly impacts the performance, as stronger solvers provide a better foundation for speedup. Since UAutomizer represents the current state-of-the-art solver, we believe future work should prioritize developing LLM-based verifiers atop state-of-the-art solvers such as UAutomizer. For a direct comparison between UAutomizer and prior LLM-based solvers, we also refer readers to Appendix A.6, where we evaluate UAutomizer on the dataset released by prior works.
Detailed Comparison with LEMUR.
To better understand the performance differences between LEMUR and Quokka, we analyze their implementation strategies and examine the instances where Quokka outperforms LEMUR. LEMUR allows LLMs to insert invariants at multiple locations (both before and at the beginning of loops) and generates multiple candidate invariants for each location. It then sequentially evaluates each candidate, verifying both the original assertion and the invariant itself. If verification times out within 30 seconds, LEMUR recursively invokes the LLM again to generate additional invariants. In contrast, Quokka adopts a much simpler strategy: it restricts invariant placement to the beginning of loops, asks the LLM to select a single location (when multiple loops exist), and generates only one invariant. The two verification queries (for the assertion and the invariant) are then executed in parallel without recursion or additional LLM invocations. Among the 353 instances that Quokka solves but LEMUR does not, 129 instances (36.5%) contain multiple loops. We hypothesize that for programs with multiple loops, LEMUR’s exhaustive search strategy becomes less effective due to the combinatorially larger space of invariant placements and candidates it must explore. In contrast, Quokka’s focused approach (i.e., requiring the LLM to commit to a single location and invariant) appears more effective in these complex scenarios. We identify three key advantages of our design. First, parallel execution of verification queries reduces wall-clock time by simultaneously checking both the assertion and the invariant. Second, constraining the LLM to select a single location and invariant forces more deliberate reasoning about which invariant will be most impactful, rather than relying on exhaustive search over many candidates. Third, eliminating recursion and repair loops reduces both implementation complexity and the risk of execution failures (which we observe in LEMUR’s logs). Together, these design choices enable Quokka to achieve superior performance while maintaining a simpler and more robust implementation.
Model # Incorrect # Goal Timeout # Invariant Timeout # Both Timeout
gpt-5.1 78 96 245 110
claude-sonnet-4 110 160 209 145
gpt-5.2 56 139 230 121
Table A7: Failure mode breakdown for the top three models on Quokka. Incorrect: the invariant is refuted by the solver. Goal Timeout: the invariant is verified, but proving the assertion under it times out. Invariant Timeout: verifying the invariant itself times out. Both Timeout: both verification queries time out.
A.11 Failure Mode Analysis
As shown in Table A7, we categorize failures into four types: 1) Incorrect Invariants: the invariant is refuted by the solver; 2) Goal Timeout: the invariant is verified, but proving the assertion under it times out; 3) Invariant Timeout: verifying the invariant itself times out; 4) Both Timeout: both verification queries time out.
A.12 Case Studies: Verification Timeout in Quokka
To better understand the verification timeouts in Quokka, we manually examined 50 timeout cases and identified five recurring categories. In this section, we select two of these categories for closer analysis and present one representative program from each. The first category includes cases where the LLM proposes a trivial but correct invariant, yet the solver still cannot complete verification within the timeout budget. The second category includes cases where the LLM proposes a correct invariant that is almost identical to the original assertion.
Case 1: Trivial and correct invariant.
Figure A2 shows a benchmark where the LLM proposes the invariant c == y after line 12. This invariant is correct because c and y are both initialized to zero before the loop and are incremented by exactly one in every iteration, so the equality is preserved throughout the loop. Moreover, it is trivial to prove: its validity follows directly from the initialization and the loop updates, and can be established by a simple inductive argument. Nevertheless, in this case, the solver still cannot prove this invariant within the timeout budget. This example shows that verification timeout may occur even when the proposed invariant is both semantically correct and trivial to prove.
1int main() {
2 short k;
3 long long y, x, c;
4 k = __VERIFIER_nondet_short();
5 assume_abort_if_not(k >= 0 && k <= 5);
6 assume_abort_if_not(k <= 256);
7
8 y = 0;
9 x = 0;
10 c = 0;
11
12 while (1) {
13 __VERIFIER_assert(-2 * y * y * y * y * y * y - 6 * y * y * y * y * y - 5 * y * y * y * y + y * y + 12 * x == 0);
14
15 if (!(c < k)) {
16 break;
17 }
18 c = c + 1;
19 y = y + 1;
20 x = y * y * y * y * y + x;
21 }
22
23 return 0;
24}
Figure A2: A timeout case where the proposed invariant c == y is semantically correct and trivial to prove, yet the solver still cannot establish it within the timeout budget.
Case 2: Correct invariant as a trivial variant of the original assertion.
A different phenomenon is shown in Figure A3, where the LLM proposes the invariant
$x=\frac{y(y+1)}{2}.$
This invariant is correct: it holds initially and is preserved by the loop updates. However, it is essentially just a trivial algebraic reformulation of the original assertion
$y^{2}-2x+y=0.$
This possibly suggests a limitation of the LLM in decomposing the verification task into a genuinely simpler intermediate property, instead of merely restating the target assertion in an equivalent form.
1int main() {
2 int k;
3 long long y, x, c;
4 k = __VERIFIER_nondet_int();
5 assume_abort_if_not(k >= 0 && k <= 10);
6
7 y = 0;
8 x = 0;
9 c = 0;
10
11 while (1) {
12
13 if (!(c < k)) {
14 break;
15 }
16
17 c = c + 1;
18 y = y + 1;
19 x = y + x;
20 }
21 __VERIFIER_assert((y * y) - 2 * x + y == 0);
22 return 0;
23}
Figure A3: A case where the proposed invariant is correct but is essentially a trivial algebraic variant of the original assertion.